OpenStudy (anonymous):

When 8 is subtracted from Ollie’s age, the result is the same as subtracting 16 from 3 times his age. How old is Ollie? ____years @freckles

OpenStudy (anonymous):

@Godlovesme @DanJS

OpenStudy (anonymous):

@ispike

OpenStudy (danjs):

pick a name for his age, let a = ollie's age

OpenStudy (danjs):

when 8 is subtracted from his age a - 8

OpenStudy (danjs):

it is the same as 16 less than 3 times his age 3a-16

OpenStudy (godlovesme):

I'll let Danjs help u :)

OpenStudy (danjs):

st those two equal, solve for a

OpenStudy (anonymous):

ok

OpenStudy (danjs):

a - 8 = 3*a - 16

OpenStudy (anonymous):

you mean 3-16?

OpenStudy (anonymous):

so a= 8?

OpenStudy (danjs):

" 16 from 3 times his age." 3 times his age - 3*a take away 16 from that 3*a - 16

OpenStudy (danjs):

solve a - 8 = 3*a - 16

OpenStudy (anonymous):

wait you mean 8+16

OpenStudy (danjs):

"When 8 is subtracted from Ollie’s age" a - 8 "the result is the same as" = "subtracting 16 from 3 times his age." 3a - 16

OpenStudy (danjs):

a - 8 = 3*a - 16 solve for a first add 16 to both sides a -8 + 16 = 3a - 16 + 16 a + 8 = 3a

OpenStudy (danjs):

take away 1a from both sides 8 = 2a divide by 2 a = 4

OpenStudy (danjs):

ollie is 4 years old right now

OpenStudy (anonymous):

Thanks